OpenStudy (anonymous):

Anyone Completed math unit 5 lesson 13 in k12??

OpenStudy (lightinganswers):

me

OpenStudy (anonymous):

what grade

OpenStudy (anonymous):

im in 6th. so i need to know

OpenStudy (anonymous):

7th im helping my little sister out with her math work

OpenStudy (anonymous):

oh. then no

OpenStudy (anonymous):

okay thanks anyway :)

OpenStudy (anonymous):

@EmogirlAtEmoooocow

OpenStudy (emogirlatemoooocow):

@thuglife187

OpenStudy (emogirlatemoooocow):

whats the question.

OpenStudy (anonymous):

Estimate by first rounding each number to the nearest integer. A. –7 B. –6 C. 6 D. 7

OpenStudy (anonymous):

\[4\frac{ 1 }{ 9 } + 2\frac{ 3 }{ 13}\]

OpenStudy (anonymous):

I did

OpenStudy (anonymous):

can you help me @Dannyreaper

OpenStudy (anonymous):

sure

OpenStudy (emogirlatemoooocow):

@thuglife187 is C . 6 1 < 5 4 1/9 rounds to 4 2 < 5 2 2/13 rounds to 2 4 + 2 = 6

OpenStudy (emogirlatemoooocow):

* its